OpenStudy (anonymous):

Ok, this is the final one of this one......... Different one........ Suppose that y is directly proportional to x. (a) Use the given formula to find the constant proportionality k. (b) Then use y=kx to find y for x=6. y=-53 when x=7 k=

OpenStudy (anonymous):

OH MY GOD!!

OpenStudy (anonymous):

Shut it Anwar......it kept changing on me! lol.

OpenStudy (anonymous):

god, all your questions are in 1 form, just the numbers that were changed. why don't you just plug the new numbers in the formula

OpenStudy (anonymous):

Because I don't get it......if I knew, I wouldn't be bothering u guys!

OpenStudy (anonymous):

a) y is directly proportional to x so y=kx when x=7, y=-53. Substitute these in y=kx so -53=k*7 k=-53/7 b) y=kx=(-53/7)x x=6 so y=(-53/7)*6=calculator

OpenStudy (anonymous):

he/she is going to post another question in the same form very soon!!

OpenStudy (anonymous):

let's bet =))

OpenStudy (anonymous):

...........

OpenStudy (anonymous):

u know what, you can just copy paste the answer, then change the number lol

OpenStudy (anonymous):

Cant do that, I have tried!

OpenStudy (anonymous):

so, you get it yet?

OpenStudy (anonymous):

WHat does y=?

OpenStudy (anonymous):

question b?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

y=_when x is 6

OpenStudy (anonymous):

y=kx, you know k and x, so just plug them in. (-53/7)*6= copy that and google? you'll get the result.

OpenStudy (anonymous):

HAHA I don't think he/she knows google

OpenStudy (anonymous):

Didn't work.

OpenStudy (anonymous):

dude, its not nice to make fun of blondie -.-

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Whatever Anwar, don't be a wingspan.

OpenStudy (anonymous):

do you have a calculator?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes I do.

OpenStudy (anonymous):

then you go - 5 3 / 7 x 6

OpenStudy (anonymous):

do you believe that she really can't do it?? I am sure that she's smarter than she seems.

OpenStudy (anonymous):

But........not in math!

OpenStudy (anonymous):

some of them don't lol. used to be one in my high school, cant even expand a(b+c)

OpenStudy (anonymous):

when are you going to post ur next question??

OpenStudy (anonymous):

I get the same equation from Google and my calculator, but it's saying the answer is incorrect.

OpenStudy (anonymous):

Well I still think she's smarter, but she is just messing with us.. if I am wrong, then I am sorry for all I have said so far.

OpenStudy (anonymous):

Correct answer was -318/7.......damn!

OpenStudy (anonymous):

whatever, nothing wrong

OpenStudy (anonymous):

(-53/7)*6=-318/7

OpenStudy (anonymous):

Too late on that one. If I do a similar exercise it will give me different numbers again.

OpenStudy (anonymous):

is you room on first floor?

OpenStudy (anonymous):

The distance D in miles to the horizon can be estimated by D (h)=1.22\[\sqrt{h}\] where h is the height of the person above the ground in feet.

OpenStudy (anonymous):

How high does a person need to be to see 26 miles?

OpenStudy (anonymous):

D=26 26=1.22sqrt(h) find h yourself. and i'm gonna put a .22 in my brain/

OpenStudy (anonymous):

No clue.

OpenStudy (amistre64):

breath in ....breath out......breath in....................uhoh, now what?

OpenStudy (anonymous):

ok, which floor are you in just now?

OpenStudy (anonymous):

2nd'

OpenStudy (anonymous):

hmmm, not high enough. Last time I throw a boy from 2nd floor and nothing happened

OpenStudy (anonymous):

What's that got to do with how many feet?

OpenStudy (anonymous):

Ok, well thanks.......I guess I am going to bed and can hopefully get these things beat asap.

OpenStudy (anonymous):

well, nothing really. I already gave you the answer up there. So are you gonna jump out of ur window? if you wont then i will

OpenStudy (anonymous):

You did?

OpenStudy (anonymous):

How? I mean don't help me if it is going to cause you to become suicidal!

OpenStudy (anonymous):

yeah, I did it. and I'm standing behind you right now hihihi

OpenStudy (anonymous):

Sigh....

OpenStudy (anonymous):

i'ts not right. 26?

OpenStudy (anonymous):

no, find square root of h first then find h. im off to bed. btw, what are u studying? Art or business or something?

OpenStudy (anonymous):

Yeah. Both.

OpenStudy (anonymous):

Sqrt of h, then find h? Wth?

OpenStudy (anonymous):

cool, but reading those things sucks. why do you have to do this math stuff anyway? Even accounting is different. D (h)=1.22√h we know D=26 so 26=1.22√h √h=26/1.22 h=(26/1.22)^2 take a sleep then start this stuff again, if u was serious about ur ability lol

OpenStudy (anonymous):

i get accounting

OpenStudy (anonymous):

accounting has nothing to do with this kind of math right?

OpenStudy (anonymous):

I still dont get feet from that

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

percents, etc, I get that.

OpenStudy (anonymous):

I bought mathway, but it doesnt seem to work for this stuff

OpenStudy (anonymous):

i got accounting last year, compulsory. hated it

OpenStudy (anonymous):

26/122^2 gets me nothing.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

i think you need a personal tutor instead

OpenStudy (anonymous):

I got two weeks and I am done with math

OpenStudy (anonymous):

(26/1.22)^2=26^2/1.22^2

OpenStudy (anonymous):

yeah, you better be without it. 2weeks huh? damn those exams

OpenStudy (anonymous):

...........

OpenStudy (anonymous):

Thank GOD...........finally!

OpenStudy (anonymous):

Go to bed..........I am. Thanks. Bad part is, I still have a lot to do!

OpenStudy (anonymous):

See I can get the equations, it's figuring out all the mumbo jumo junk

OpenStudy (anonymous):

not the worst thing in the world lol. u should be happy. I have to stick to maths my whole life. fml.

OpenStudy (anonymous):

Hey, u must like it!

OpenStudy (anonymous):

no way i like it haha. i like those machinery stuff but not maths, which always comes along. I'm here just because fb anf utube waste me too much time. So I rather waste my time on helping people abit heer.

OpenStudy (anonymous):

Good deal!

OpenStudy (anonymous):

anyway. im off to bed too. the birds are chirping already lol. goognight

OpenStudy (anonymous):

morning for me though